OpenStudy (skullpatrol):

Find a number between 0.999... and 1.

OpenStudy (anonymous):

lim x->1-

OpenStudy (anonymous):

0.9999

OpenStudy (anonymous):

Is this a trick question? :/

OpenStudy (anonymous):

-0.999

OpenStudy (skullpatrol):

It is a "make you think" question.

OpenStudy (skullpatrol):

Does there exist an x such that $$\Huge0.999...<x<1$$

OpenStudy (skullpatrol):

x must be larger than 0.999... AND smaller than 1

OpenStudy (anonymous):

there are infinite number between the open interval \((0.999, 1)\)

OpenStudy (rational):

\[0.999\ldots = \sum\limits_{n=1}^{\infty} \dfrac{9}{10^n} = 1\]

OpenStudy (rational):

question is same as finding a number between 1 and 1

OpenStudy (mathmath333):

so its empty set

OpenStudy (skullpatrol):

If you cannot find a number between 0.999... and 1, then they are the same number.

OpenStudy (rational):

empty set is a correct answer

OpenStudy (skullpatrol):

The objective is to write down the decimal number which is larger than 0.999... but less than 1.

OpenStudy (anonymous):

@skullpatrol People have been debated that for years now. Whether or not 0.999 equals 1. Some say it does, while others say it does not.

OpenStudy (skullpatrol):

Try to "think about it" @Night-Watcher How can you make 0.999... any bigger?

OpenStudy (rational):

0.999 is not 1 0.999... is exactly same as 1 however

OpenStudy (anonymous):

By adding other 9's at the end.

OpenStudy (skullpatrol):

But there are infinitely many 9s on the end...it is endLESS

OpenStudy (skullpatrol):

Just like 1/3 = 0.333...

OpenStudy (skullpatrol):

but see 2/3 = 0.666...

OpenStudy (skullpatrol):

1/3 + 2/3 = ?

OpenStudy (anonymous):

1

OpenStudy (skullpatrol):

0.333... + 0.666... = ?

OpenStudy (anonymous):

0.999....

OpenStudy (mathmath333):

there are many tricks \(\large \begin{align} \color{black}{x=0.99999999......\hspace{.33em}\\~\\ 10x=9.9999999......\hspace{.33em}\\~\\ 10x=9+0.9999999......\hspace{.33em}\\~\\ 10x=9+x\hspace{.33em}\\~\\ 9x=9\hspace{.33em}\\~\\ x=1 }\end{align}\)

OpenStudy (anonymous):

It's equal to 1, so there can't be a number between. 0.999 as a decimal equals 9/9 which equals 1!

OpenStudy (anonymous):

0.999000 can be written as the fraction 1

OpenStudy (mathmath333):

0.999000 it is different from 0.99999........

OpenStudy (anonymous):

10x = 9.99 - x = 0.999 ---------- 9x = 8.991.

OpenStudy (anonymous):

to solve 9x=9 we divide both sides of it by 9, and we'll get that x=9/9.

OpenStudy (anonymous):

Then 9/9 reduced is 1

OpenStudy (skullpatrol):

@Night-Watcher you're missing the idea of "infinitely many" nines...

OpenStudy (anonymous):

*sigh* What i'm going to be doing after finding out the answer. http://wac.450f.edgecastcdn.net/80450F/thefw.com/files/2013/05/SPoiler.gif

OpenStudy (skullpatrol):

If I want to increase the size of a decimal number I must increase one of its digits to the next larger one. For example, 0.752 < 0.762 because I increased 5 to 6, right?

OpenStudy (skullpatrol):

Or 0.333... < 0.343... because I increased 3 to 4.

OpenStudy (skullpatrol):

Now with 0.999... if I increase any of the 9s by one I will get a 0, For example 0.999... + 0.01 = 1.009...

OpenStudy (skullpatrol):

But look that is larger than 1.

OpenStudy (anonymous):

0.999 + 0.001= 1

OpenStudy (anonymous):

And 0.999 + 0.0001= 0.9991

OpenStudy (anonymous):

So would that number be between the two?

OpenStudy (skullpatrol):

Nope. You need to add 0.999... + 0.0001

OpenStudy (skullpatrol):

There is no number 0.0001 that you can add to 0.999... that will not make it larger than 1

OpenStudy (skullpatrol):

00.9999999... +0.00001 ----------- 1.00000999... > 1

OpenStudy (skullpatrol):

Any tiny, tiny number you add to 0.999... will make it greater than 1, so that must mean there is no number "between" and that can only be if they are the same

OpenStudy (anonymous):

Touché.

OpenStudy (anonymous):

If we have 0.000000000000................................................1 at the last, then??

OpenStudy (xapproachesinfinity):

the empty set makes a good answer like @mathmath333 said

OpenStudy (skullpatrol):

@eta if you have a "last 1" on the end of your "0.000000000000................................................1." Then 0.9999999999999999999999999999999999999999... +.000000000000................................................1 ---------------------------------------------------- 1.0000000000000000000000000000000000000999... > 1

OpenStudy (anonymous):

At the end means, where 9 will end.. If it is not ending, then add that 1 to that non-ending end. :P

OpenStudy (skullpatrol):

Can there be an end to something that does not end? @eta

OpenStudy (skullpatrol):